Express the given number in a normal form: (2 x 1000) + (2 x 10)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(2 x 1000) + (2 x 10) will be written in a normal form in the following way:

(2 x 1000) + (2 x 10) = 2000 + 20 = 2020

Thanks